OpenStudy (anonymous):

√(5)^(8n)=125^(n+3) solve for n

OpenStudy (anonymous):

square both sides first

OpenStudy (anonymous):

you get 5^(16n) = 125^(2n+6)

OpenStudy (anonymous):

5^(16n) = 5^3(2n+6)

OpenStudy (anonymous):

16n = 6n + 18 16n-6n = 18 10n = 18 n =9/5

OpenStudy (anonymous):

thanks